OpenStudy (anonymous):

How to integrate exp(x) / (1-exp(x) )^2 please help

OpenStudy (anonymous):

Use a u substitution for \(u = 1-e^x \implies \frac{du}{-e^x} = dx\)

OpenStudy (anonymous):

Then integrate with respect to u.

OpenStudy (anonymous):

Thanks a lot I wil try it and let u know

OpenStudy (anonymous):

yes it working thank you polpak